OpenStudy (anonymous):

Complex Rational Expression 1/a + 1/b --------- 1/a - 1/b Medal will be awarded!!!!

OpenStudy (isaiah.feynman):

\[\frac{ \frac{ 1 }{ a }+\frac{ 1 }{ b } }{ \frac{ 1 }{ a }-\frac{ 1 }{ b } } = \frac{ \frac{ b+a }{ ab } }{ \frac{ b-a }{ ab } }= (b-a)(b+a)=b^{2}-a^{2} \]

OpenStudy (anonymous):

how come you multiplied by ab?

OpenStudy (anonymous):

wait what did you multiply by for the common denominators?

OpenStudy (isaiah.feynman):

I used this formula \[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad -bc}{ bd }\]

OpenStudy (isaiah.feynman):

I mean "+" on the right side.

OpenStudy (isaiah.feynman):

Not minus